OpenStudy (anonymous):

what does the m in 4+m/8=3/4 stand for

OpenStudy (anonymous):

\[\frac{4+m}{8}=\frac{3}{4}\]

OpenStudy (anonymous):

4+ m/8 = 3/4 m/8 = 3/4 - 4 m/8 = (3-16)/4 m/8 = -13/4 m = -26/4

OpenStudy (anonymous):

if it is (4+m)/8 = 3/4 4+m = (3)(8)/4 4+m = 6 m = 2

OpenStudy (anonymous):

It could be any letter,not just m but a or b, etc, it is what is called a variable, a stand-in for a number that you will calculate as has been done above...

Parth (parthkohli):

Here, \(\rm m\) is your unknown constant.